# Are We Truly Forgetting? A Critical Re-examination of Machine Unlearning Evaluation Protocols

### v2

@@ B.1 Training and Retraining Details
For the experiments, both the original and retrained models are required. In the case of ResNet-50, we trained it from scratch. For ConvNeXt-T and Swin-T, we initialized them with ImageNet-21K pretrained weights [19]. Detailed configuration information is provided in Table 2.
Table 2: Training Details for Original and Retrain Models.
@@
B.2 Unlearning Details
We refer to the hyperparameters from the original studies. However, because most prior works proposed and tested these methods primarily on smaller datasets (e.g., CIFAR-10, CIFAR-100), critical implementation details are not fully specified when removing multiple classes from large-scale datasets. To address this limitation, we perform a grid search for each method to identify the optimal learning rate and number of epochs. Specifically, we retrain the model (excluding the forget set $\mathcal{D}_{f}$) using the same learning rate and epochs as in the original training for the Retrain method. We limit FT at 40 epochs to mitigate excessive time costs. For PL, GA, RL, and SalUn, we find that strong unlearning can be achieved within 10 epochs on large-scale data. Finally, considering the balance of logit-based and representation-based metrics, we fix the maximum epochs for DUCK, CU, SCAR, and SCRUB at 15, 80, 20, and 90, respectively. These upper limits are determined first under the scenario of Random 100 class unlearning. For Random $k$ class experiments ($k\in\{100,200,300\}$), we choose the actual number of epochs while respecting these maximum values, guided by the H-LR score. For Top $k$ class unlearning, we reuse the hyperparameter configurations established in the corresponding Random $k$ setting. All experiments were conducted using a single NVIDIA RTX 4090 GPU.
Appendix C Additional Experiments
We also provide t-SNE visualization of all models in Fig. 8. Our full results in all unlearning baselines across all unlearning scenarios can be found in Tables 456, and 7.
Table 3: Unlearning evaluation with 100 random class-wise forgetting on ConNeXt and Swin Transformers.
ImageNet-1K Office-Home, CUB, DomainNet-126

### v3

@@ Appendix A Implementation Details
For the experiments, both the original and retrained models are required. In the case of ResNet-50, we trained it from scratch. For ConvNeXt-T and Swin-T, we initialized them with ImageNet-21K pre-trained weights, following [22]. Detailed configuration information is provided in Table 6.
@@
Table 6: Training details for original and retrain models.
Unlearning details. We refer to the hyperparameters from the original studies. However, because most prior works proposed and tested these methods on smaller datasets (e.g., CIFAR-10, CIFAR-100), critical implementation details are not specified when removing multiple classes from large-scale datasets. To cope with this, we perform a grid search for each method to identify the optimal learning rate and the number of epochs. Specifically, we retrain the model (excluding the forget set) using the same learning rate and number of epochs as in the original training for the Retrain method. We limit FT to 40 epochs to mitigate excessive time costs. For PL, GA, RL, and SalUn, we find that strong unlearning can be achieved within 10 epochs on large-scale data. Finally, considering the balance of logit-based and representation-based metrics, we fix the maximum epochs for DUCK, CU, SCAR, and SCRUB at 15, 80, 20, and 90, respectively. These upper limits are first determined under the scenario of random 100 class unlearning. For Random-N class experiments ($N\in\{100,200,300\}$), we choose the actual number of epochs while respecting these maximum values and are guided by the H-LR score. For Top-N class unlearning, we reuse the hyperparameter configurations established in the corresponding Random-N setting. All experiments were conducted using a single NVIDIA RTX 4090 GPU.
Appendix B Additional Experimental Results
Our comprehensive results in all unlearning baselines across all unlearning scenarios can be found in Tables 9-15.
Different backbone architectures. Table 7 provides an extended comparison of the baselines using alternative backbones, specifically ConvNeXt-T [32] and Swin-T [31]. While our primary analysis was conducted using ResNet-50 (as it is the standard backbone in the unlearning literature), we included these modern architectures to confirm that our benchmark’s findings are robust and not tied to a specific architectural design. As the results in Table 7 demonstrate, the similar trends observed with ResNet-50 re-emerge. Notably, the limitations of certain methods (e.g., GA, RL, and SalUn showing poor representation-level unlearning) and the relative strengths of methods like PL and CU hold across these different backbones. This consistency confirms that our framework is capable of reliably assessing unlearning effectiveness across diverse architectures, demonstrating that the identified issues (e.g., the gap between logit-based and representation-based forgetting) are fundamental to the unlearning algorithms themselves and not an artifact of a single backbone. We compare only model-agnostic methods.
ImageNet-1K Office-Home, CUB, DomainNet